# A simplified model for coupling Darrieus–Landau and diffusive-thermal instabilities

### 24 Mar 2026 v1

@@ Abstract
Darrieus–Landau instability , Diffusive-thermal instability , Michelson–Sivashinsky equation , Kuramoto–Sivashinsky equation
@@ 1 Introduction
The theory of instability in premixed flames has, for the most part, evolved along two parallel and only partially intersecting lines of inquiry. The first of these, originating in the classical analyses of Darrieus [darrieus1938propagation] and Landau [landau1944slow], establishes that a plane flame, by virtue of the density change induced through thermal expansion, is intrinsically susceptible to a hydrodynamic instability of long wavelength. The second instability, foreseen by Zeldovich [zeldovich1944theory] in the same era, arises from the disparity between thermal and molecular diffusion, and bears a formal resemblance to the Turing instability. The rigorous mathematical development of the Darrieus–Landau (DL) and diffusive-thermal (DT) instabilities saw a significant revival in the late 1970s and early 1980s through the pioneering contributions of Sivashinsky, Clavin, Matalon, Joulin, and their collaborators. Even today, these instabilities are largely treated separately rather than within a unified framework. While the DL instability persists in premixed flames under general conditions, the DT instability arises only for sub-unity Lewis numbers, due to preferential transport of heat relative to species diffusion [williams2018combustion, clavin2016combustion].
@@
The Matalon–Matkowsky–Clavin–Joulin theory  [clavin1982effects, matalon1982flames, clavin1983premixed, pelce1988influence] has provided a definitive account of the DL instability, revealing how the inner structure of the flame imparts a stabilising influence at short wavelengths, of the order of the flame thickness. In a parallel development, a quantitative theory of the diffusive-thermal instability was formulated by Sivashinsky, Clavin, and Joulin [clavin1985effect, sivashinsky1977diffusional], but in this treatment the full hydrodynamic effects were suppressed, an approximation (i.e., thermo-diffusive approximation) that eliminates the DL mechanism. A comprehensive treatment in which both mechanisms emerge naturally from the full set of conservation equations remains, even now, an undertaking of considerable difficulty. It is of some historical interest that Sivashinsky, in his seminal contribution of 1977 [sivashinsky1988nonlinear] (see also the review of Matalon [matalon2007intrinsic]), had already envisaged a regime in which the two instabilities might be coupled, though of a character distinct from that considered here.
@@
In addition to theoretical investigations, a substantial body of numerical and experimental work has been devoted to the study of these instabilities, both singly and in combination; see, for example, [jackson1984effect, denet1992numerical, denet1995numerical]. In light of the complexity of a complete theory, it is desirable to construct a simplified phenomenological model that couples these instabilities. Such a model can provide insight into the interaction between DL and DT modes without requiring the full, detailed analysis. The purpose of the present note is to advance such a model. Crucially, while previous models often treat these instabilities as additive, we identify a ’hydro-diffusive’ coupling that becomes dominant near the stability threshold. The objective of this short note is precisely to present such a minimal coupled model, introducing a new physical parameter, the hydro-diffusive area, to characterise this interaction.
@@ 2 A modified dispersion relation near the instability onset (weak heat release)
A simple combination of these two linear relations might suggest [sivashinsky1988nonlinear, matalon2007intrinsic]
@@ 3 Weakly nonlinear evolution: Two regimes
The instability onset threshold is dictated by the limit $\varepsilon\to 0^{+}$. Depending on how $\mathcal{M}$ scales with $\varepsilon$, we can consider two distinct regimes where a weakly nonlinear description can be provided. Our objective is here to write down the governing equations for the flame amplitude $f(x,t)$, i.e., departure of flame excursion from the planar state, following the approach in [rajamanickam2024tricritical, daou2024diffusive].
@@ 3.1 The classical Michelson–Sivashinsky equation: $\mathcal{M}\gg\sqrt{\varepsilon\mathcal{N}}$
in which the scaling law for $f\sim\sigma/k^{2}$ corresponds to the weakly nonlinear regime, i.e., $f_{t}\sim\tfrac{1}{2}f_{x}^{2}$. In this case, one covers the classical Michelson–Sivashinsky equation [clavin2016combustion]
@@
3.2 The DL-DT crossover regime: $\mathcal{M}\sim\sqrt{\varepsilon\mathcal{N}}$
@@
3.3 Phenomenological approach to the regime where $\mathcal{M}\ll-\sqrt{\varepsilon\mathcal{N}}$
The regime $\mathcal{M}\ll-\sqrt{\varepsilon\mathcal{N}}$ lies deep within the unstable region where both DL and DT instabilities are strongly active. In this regime, no single asymptotic scaling exists that would yield a weakly nonlinear reduction — the linear growth rate $\sigma=a|k|-|\mathcal{M}|k^{2}-\mathcal{N}|k|^{3}$ involves multiple competing effects at $O(1)$ wavenumbers, and a systematic asymptotic expansion in the small parameter $\varepsilon$ is no longer possible. Nevertheless, one may attempt a heuristic phenomenological model that bridges the Michelson–Sivashinsky equation (describing the DL-dominated regime) and the Kuramoto–Sivashinsky equation (describing the DT-dominated regime). The idea is to retain all terms that appear in the linear dispersion relation, including the ultimate short-wavelength cutoff provided by the flame thickness, which would enter as a $4k^{4}$ term in Fourier space. A minimal phenomenological extension that captures the combined effects is
@@ 3.3 Phenomenological approach to the regime where $\mathcal{M}\ll-\sqrt{\varepsilon\mathcal{N}}$
where $\mathcal{M},\mathcal{N}\in\mathbb{R}$. Once again, we emphasise that this equation is not derived asymptotically but rather constructed heuristically. While lacking asymptotic justification, such a phenomenological model may nonetheless capture qualitative features of the strongly nonlinear interaction between Darrieus–Landau and diffusive-thermal instabilities when both are significant.
@@ 4 Illustrative numerical results
The domain is taken to be periodic, of length $2\pi L$; in dimensional units this corresponds to a length of order $L\times\delta_{L}/\sqrt{\varepsilon}$. From a numerical standpoint, it is convenient to to work with the slope variable $G=F_{\xi}$; in this formulation the gradual drift of the mean value of $F$ is avoided [rajamanickam2024hydrodynamic]. Furthermore, the computational domain may be held fixed by introducing the variables $X=\xi/L$ and $T=\tau/L$ such that
@@
where $\mu=\mathcal{M}/\sqrt{\varepsilon}\sim O(1)$ provides a measure of the Markstein number and $\nu=1/L^{2}>0$ is is inversely proportional to the square of the domain size. The initial and boundary conditions are given by
@@
For even smaller domains, the $\mu=+1$ case does not yield further interesting dynamics for $\nu$ of the order $10^{-3}$ or greater, whereas the $\mu=-1$ case continues to exhibit rich behaviour, as illustrated for three values of $\nu$ in Fig. 3. To summarise, the new proposed equation (12) exhibits fundamentally different characteristics depending on the sign of the Markstein number. For $\mathcal{M}>0$ (i.e., $\mu>0$), the dynamics resembles that of the Michelson–-Sivashinsky equation, though it may not now be represented by a finite number of poles and cusp coalescence; this is true for moderately large domains, because in very large domains, chaos appears even when $\mathcal{M}>0$, as demonstrated in Fig. 4. For $\mathcal{M}<0$ (i.e., $\mu<0$), the behaviour is akin to that of the Kuramoto–-Sivashinsky equation, corresponding to an effectively infinite-dimensional dynamical system with sustained chaotic activity.
@@
In sufficiently large domains, regardless of the sign of $\lambda$, two competing effects are constantly at play: the tendency to form large-scale DL cusp structures and the propensity to generate small-scale cellular wrinkles. These two tendencies perpetually interact, each influencing and often overtaking the other, either sequentially or simultaneously. The resulting dynamics is thus a ceaseless competition between coherent structure formation and fine-scale disruption, giving rise to the complex, recurrent cycles observed in the simulations. The present model thus provides a unified description spanning both regimes, bridging the gap between two classical paradigms of flame front instability while also revealing this richer, emergent behaviour. The present model thus provides a unified description spanning both regimes, bridging the gap between two classical paradigms of flame front instability.
@@ 5 Concluding remarks
Figure 5: Marginal stability curves for different values of $\epsilon=(r-1)/2$ in the limit $\beta\to\infty$ where $\beta$ is the Zeldovich number. The red dashed line corresponds to the formula $l=-2-8k^{2}$, applicable when $r=1$ (i.e., $\epsilon=0$), obtained by Sivashinsky [sivashinsky1977diffusional]. The four solid lines correspond to $r=\{1.2,1.5,2,5\}$ (i.e., $\epsilon=\{0.1,0.25,0.5,2.0\}$), obtained by Jackson and Kapila [jackson1984effect]. The green shaded region indicates the DL–DT crossover regime where the cubic term $\mathcal{A}|k|^{3}$ identified in the present work becomes significant; the red shaded region indicates where it is negligible.
@@
It is perhaps to be expected that the familiar pole decomposition, which so elegantly characterises the Michelson–Sivashinsky equation, does not extend to this equation. The $\tfrac{3}{2}$-Laplacian, $(-\Delta)^{3/2}f$, fundamentally alters the singularity structure of the solutions, precluding a finite-pole representation. This suggests that the dynamics encompassed by the present equation may be considerably richer than those of the pure DL case, in which the asymptotic state is typically governed by the coalescence of poles into a single cusp. This mathematical complexity mirrors the physical intricacy of the regime under consideration. By incorporating the coupling between hydrodynamic and diffusive-thermal effects, the proposed model offers a more robust framework for interpreting the fine-scale cellular structures that have been observed in experiments on flames near the threshold of stability. Within this distinguished crossover limit, the hydro-diffusive area $\mathcal{A}$ serves as a robust coupling constant. Whereas the Markstein number $\mathcal{M}$ is highly sensitive to the Lewis number, changing sign as the DT instability threshold is crossed, the coefficient $\mathcal{N}$ remains a stable, leading-order interaction term. To a first approximation, $\mathcal{N}$ may be treated as independent of the Lewis number in this narrow regime; however, it may be expected to vary as the internal flame structure undergoes more substantial modifications farther from the threshold. A first-principles derivation of the coefficient $\mathcal{N}$ from the full Navier–Stokes equations remains a challenging undertaking, requiring an asymptotic development that goes beyond the established Matalon–Matkowsky–Clavin–Joulin theory. Its determination may prove more tractable within the framework of Darcy’s law, where the hydrodynamic description is considerably simplified [rajamanickam2024hydrodynamic, rajamanickam2026flame]. Alternatively, $\mathcal{N}$ might be inferred from experimental measurements of cellular flame structures or extracted from direct numerical simulations, thereby providing a practical route to quantifying the hydro-diffusive area $\mathcal{A}$. More specifically, the limit $\mathcal{M}\to 0^{\pm}$ is of particular interest; in this limit, the classical theory, with its assumption of an order-unity Markstein number, becomes inapplicable. A new hydrodynamic description of premixed flames is therefore required, a direction that is both warranted and reserved for future investigation.
@@
Note that the present study has been conducted within the framework of a weak-heat-release approximation, adopted in the interest of simplicity. The qualitative features of the discussion may be expected to carry over, with appropriate modifications, when this approximation is relaxed. In the latter case, for instance, the quadratic stabilisation vanishes not simply when $\mathcal{M}\leq 0$, but rather when $\mathcal{M}-\mathcal{M}_{c}\leq 0$, where $\mathcal{M}_{c}$ denotes a critical Markstein number influenced by effects such as viscosity (see, for example, formula (19) and $\S$3.2 in [matalon2018darrieus], which follows from a Taylor-series expansion of the Clavin–Garcia equation).
@@
Finally, it is appropriate to conclude this note by comparing (see Fig. 6) the distinguished DL-DT crossover regime proposed by Sivashinsky [sivashinsky1988nonlinear, matalon2007intrinsic] for small values of $\mathcal{M}$ (or $\mathcal{M}-\mathcal{M}_{c}$) with ours:
@@
which complements the pure DL regime scalings:
@@
Since our Markstein-number scaling $\mathcal{M}\sim\sqrt{\varepsilon}$ characteristic of the present model exceeds that of Sivashinsky $\mathcal{M}\sim\varepsilon^{2/3}$, the cubic stabilisation becomes operative earlier as the Markstein number is reduced toward the DT instability threshold. In other words, as a flame transitions from DL-modulated propagation toward the DT instability threshold, the cubic term engages first, providing the primary stabilising mechanism, while the quartic term only emerges in the deeper asymptotic limit where $\mathcal{M}$ is extremely small. Accordingly, when comparing cubic $(-|k|^{3})$ versus quartic $(-k^{4})$ stabilisation, the cubic model yields cellular structures of larger characteristic dimension, slower growth rates, and greater amplitudes. The present scaling thus captures the physically more immediate crossover regime where both hydrodynamic and diffusive-thermal instabilities participate on an equal footing.

### 07 Apr 2026 v2

@@ Abstract
Darrieus–Landau instability , Diffusive-thermal instability , Michelson–Sivashinsky equation , Kuramoto–Sivashinsky equation , Weakly nonlinear analysis , DL-DT crossover regime
@@ 1 Introduction
The theory of instability in premixed flames has, for the most part, evolved along two parallel and only partially intersecting lines of inquiry. The first of these, originating in the classical analyses of Darrieus darrieus1938propagation and Landau landau1944slow, establishes that a plane flame, by virtue of the density change induced through thermal expansion, is intrinsically susceptible to a hydrodynamic instability of long wavelength. The second instability, foreseen by Zeldovich zeldovich1944theory in the same era, arises from the disparity between thermal and molecular diffusion, and bears a formal resemblance to the Turing instability. The rigorous mathematical development of the Darrieus–Landau (DL) and diffusive-thermal (DT) instabilities saw a significant revival in the late 1970s and early 1980s through the pioneering contributions of Sivashinsky, Clavin, Matalon, Joulin, and their collaborators. Even today, these instabilities are largely treated separately rather than within a unified framework. While the DL instability persists in premixed flames under general conditions, the DT instability arises only for sub-unity Lewis numbers, due to preferential transport of heat relative to species diffusion williams2018combustion, clavin2016combustion.
@@
The Matalon–Matkowsky–Clavin–Joulin (MMCJ) theory  clavin1982effects, matalon1982flames, clavin1983premixed, pelce1988influence has provided a definitive account of the DL instability, revealing how the inner structure of the flame imparts a stabilising influence at short wavelengths, of the order of the flame thickness. In a parallel development, a quantitative theory of the diffusive-thermal instability was formulated by Sivashinsky, Clavin, and Joulin clavin1985effect, sivashinsky1977diffusional, but in this treatment the full hydrodynamic effects were suppressed, an approximation (i.e., thermo-diffusive approximation) that eliminates the DL mechanism. A comprehensive treatment in which both mechanisms emerge naturally from the full set of conservation equations remains, even now, an undertaking of considerable difficulty. It is of some historical interest that Sivashinsky, in his seminal contribution of 1977 sivashinsky1988nonlinear (see also the review of Matalon matalon2007intrinsic), had already envisaged a regime in which the two instabilities might be coupled, though of a character distinct from that considered here.
@@
In addition to theoretical investigations, a substantial body of numerical and experimental work has been devoted to the study of these instabilities, both singly and in combination; see, for example, jackson1984effect, denet1992numerical, denet1995numerical. In light of the complexity of a complete theory, it is desirable to construct a simplified phenomenological model that couples these instabilities. Such a model can provide insight into the interaction between DL and DT modes without requiring the full, detailed analysis. The purpose of the present note is to advance such a model. Crucially, while previous models often treat these instabilities as additive, we identify a ’hydro-diffusive’ coupling that becomes dominant near the stability threshold. The objective of this short note is precisely to present such a minimal coupled model, introducing a new physical parameter, the hydro-diffusive area, to characterise this interaction.
@@ 2 A modified dispersion relation near the instability onset (weak heat release)
A simple combination of these two linear relations might suggest sivashinsky1988nonlinear, matalon2007intrinsic
@@ 3 Weakly nonlinear evolution: Two regimes
The instability onset threshold is dictated by the limit $\varepsilon\to 0^{+}$. Depending on how $\mathcal{M}$ scales with $\varepsilon$, we can consider two distinct regimes where a weakly nonlinear description can be provided. Our objective is here to write down the governing equations for the flame amplitude $f(x,t)$, i.e., departure of flame excursion from the planar state, following the approach in rajamanickam2024tricritical, daou2024diffusive.
@@ 3.1 The classical Michelson–Sivashinsky equation: $\mathcal{M}\gg\sqrt{\varepsilon\mathcal{N}}$
in which the scaling law for $f\sim\sigma/k^{2}$ corresponds to the weakly nonlinear regime, i.e., $f_{t}\sim\tfrac{1}{2}f_{x}^{2}$. In this case, one covers the classical Michelson–Sivashinsky equation clavin2016combustion
@@
3.2 The DL-DT crossover regime: $\mathcal{M}\sim\sqrt{\varepsilon\mathcal{N}}$ with $\mathcal{N}\sim O(1)$
@@
3.3 Heuristic approach to the crossover regime, $\mathcal{M}\sim\sqrt{\varepsilon\mathcal{N}}$, when $\mathcal{N}$ itself is small
The crossover regime when $\mathcal{N}$ is a small number lies in the deep asymptotic limit $\mathcal{M}\to 0$. One may attempt a heuristic phenomenological model that bridges the Michelson–Sivashinsky equation (describing the DL-dominated regime) and the Kuramoto–Sivashinsky equation (describing the DT-dominated regime). The idea is to retain all terms that appear in the linear dispersion relation, including the ultimate short-wavelength cutoff provided by the flame thickness, which would enter as a $4k^{4}$ term in Fourier space. A minimal phenomenological extension that captures the combined effects is
@@
This equation, which pertains to the linear dispersion relation $\sigma=\varepsilon|k|-\mathcal{M}k^{2}-\mathcal{N}|k|^{3}-4k^{4}$, implies the scaling laws
$k\sim\varepsilon^{1/3},\qquad\sigma\sim\varepsilon^{4/3},\qquad f\sim\varepsilon^{2/3},\qquad\mathcal{M}\sim\varepsilon^{2/3},\qquad\mathcal{N}\sim\varepsilon^{1/3}.$ (14)
That is to say, both $\mathcal{M}$ and $\mathcal{N}$ must be small numbers, although $\mathcal{N}\gg\mathcal{M}$. Such a scenario, in which both coefficients vanish simultaneously with specific scaling exponents, is a special (codimension-two) coincidence in parameter space. Generically, as the DT threshold is approached ($\mathcal{M}\to 0$), the hydro-diffusive number $\mathcal{N}$ is expected to remain order unity unless a separate fine-tuning occurs. Therefore, the deep asymptotic limit described here is unlikely to be realised in practice; the generic crossover regime is that discussed in Section 3.2.
@@ 4 Illustrative numerical results
The domain is taken to be periodic, of length $2\pi L$; in dimensional units this corresponds to a length of order $L\times\delta_{L}/\sqrt{\varepsilon}$. From a numerical standpoint, it is convenient to to work with the slope variable $G=F_{\xi}$; in this formulation the gradual drift of the mean value of $F$ is avoided rajamanickam2024hydrodynamic. Furthermore, the computational domain may be held fixed by introducing the variables $X=\xi/L$ and $T=\tau/L$ such that
@@
where $\mu=\mathcal{M}/\sqrt{\varepsilon}\sim O(1)$ provides a measure of the Markstein number and $\nu=1/L^{2}>0$ is inversely proportional to the square of the domain size. The initial and boundary conditions are given by
@@
For even smaller domains, the $\mu=+1$ case does not yield further interesting dynamics, provided $\nu\gtrapprox 10^{-3}$, whereas the $\mu=-1$ case continues to exhibit rich dynamic behaviour, as illustrated for three values of $\nu\sim 10^{-2}$ in Fig. 3. To summarise, the new proposed equation (12) exhibits fundamentally different characteristics depending on the sign of the Markstein number. For $\mathcal{M}>0$ (i.e., $\mu>0$), the dynamics resembles that of the Michelson–-Sivashinsky equation, though it may not now be represented by a finite number of poles and cusp coalescence; this is true for moderately large domains, because in very large domains, chaos appears even when $\mathcal{M}>0$, as demonstrated in Fig. 4. The last observation may perhaps have some relevance to the fractal structures reported in yu2015fractal for pure DL instability. For $\mathcal{M}<0$ (i.e., $\mu<0$), the behaviour is akin to that of the Kuramoto–-Sivashinsky equation, corresponding to an effectively infinite-dimensional dynamical system with sustained chaotic activity.
@@
In sufficiently large domains, regardless of the sign of $\lambda$, two competing effects are constantly at play: the tendency to form large-scale DL cusp structures and the propensity to generate small-scale cellular wrinkles. These two tendencies perpetually interact, each influencing and often overtaking the other, either sequentially or simultaneously. The resulting dynamics is thus a ceaseless competition between coherent structure formation and fine-scale disruption, giving rise to the complex, recurrent cycles observed in the simulations. The present model thus provides a unified description spanning both regimes, bridging the gap between two classical paradigms of flame front instability while also revealing this richer, emergent behaviour.
5 Can the cubic term be destabilising?
After this brief note was written, the author was made aware of the latest publication by Bechtold and Matalon bechtold2026long, which similarly explores a coupled model emphasizing a cubic term. Their derivation, based on the standard multi-scale theory, reveals an unexpected sensitivity to the Prandtl number (Fig. 3 in bechtold2026long). Specifically, for realistic Prandtl numbers (e.g., $\textit{Pr}=0.7$), their cubic correction is destabilising for all Lewis numbers of practical interest, including at $Le=1$, indicating a new type of instability on top of DL and DT. A similar implication was hinted at in the earlier work of Class et al. class2003unified, who alluded to a negative surface-tension-like effect within the same multi-scale framework. While these derivations represent valuable steps toward a unified description, it is important to recognise that the MMCJ theory is fundamentally a long-wave expansion, formally valid only for small wavenumbers, with the small parameter being the ratio of flame thickness to hydrodynamic scale. Extending this expansion to cubic terms may not be rigorously justified; the theory’s limit may have already been reached at the quadratic level, and cubic-order predictions, whether stabilising or destabilising, should be interpreted with caution. Thus, the prediction of a cubic destabilisation suggests that either there is a new instability lurking, which has not been delineated from DL in experiments and simulations, or a careful re-examination of the classical multi-scale framework is required, especially in the DL-DT crossover regime. A non-standard asymptotic analysis may be needed to resolve the sign of this term.
From the phenomenological perspective adopted here, if a cubic term is taken to be destabilising, the system’s saturation would necessarily rely on higher-order quartic terms, leading to the scaling laws discussed in Section 3.3 in the weak heat-release approximation. This requires a specific (codimension-two) alignment where $\mathcal{M}$ and $\mathcal{N}$ vanish simultaneously, i.e.,
$\mathcal{N}=\mathcal{N}(\mathcal{M})\quad\text{such that}\quad\mathcal{N}(\mathcal{M}\to 0)\to 0\quad\text{when}\quad\varepsilon\ll 1.$ (20)
In the present framework, however, the cubic term is interpreted as a direct hydro-diffusive coupling that persists independently of the product $\epsilon\mathcal{M}$. By defining the coupling constant as $(C_{0}+C_{1}\epsilon\mathcal{M})|k|^{3}$, the present model targets the regime where $C_{0}\neq 0$ is dominant, thereby providing a robust stabilising mechanism. This stands in contrast to the case where $C_{0}\ll 1$, in which both the cubic and quadratic terms vanish together, requiring a quartic term ($4k^{4}$) for stabilisation. The present work advocates for $C_{0}\neq 0$ as the physically immediate crossover regime, representing a strong, non-perturbative coupling between hydrodynamics and the internal flame structure. This interaction is potentially invisible to standard perturbative asymptotics, which treat the ratio of flame thickness to hydrodynamic length as a small parameter. The possibility $C_{0}\ll 1$, and the associated deep asymptotic limit of Section 3.3, is thus left as a restricted, fine-tuned (codimension-two) special case.
@@ 6 Concluding remarks
Figure 5: Marginal stability curves for different values of $\epsilon=(r-1)/2$ in the limit $\beta\to\infty$ where $\beta$ is the Zeldovich number. The red dashed line corresponds to the formula $l=-2-8k^{2}$, applicable when $r=1$ (i.e., $\epsilon=0$), obtained by Sivashinsky sivashinsky1977diffusional. The four solid lines correspond to $r=\{1.2,1.5,2,5\}$ (i.e., $\epsilon=\{0.1,0.25,0.5,2.0\}$), obtained by Jackson and Kapila jackson1984effect. The green shaded region indicates the DL–DT crossover regime where the cubic term $\mathcal{A}|k|^{3}$ identified in the present work becomes significant; the red shaded region indicates where it is negligible.
@@
It is perhaps to be expected that the familiar pole decomposition, which so elegantly characterises the Michelson–Sivashinsky equation, does not extend to this equation. The $\tfrac{3}{2}$-Laplacian, $(-\Delta)^{3/2}f$, fundamentally alters the singularity structure of the solutions, precluding a finite-pole representation. This suggests that the dynamics encompassed by the present equation may be considerably richer than those of the pure DL case, in which the asymptotic state is typically governed by the coalescence of poles into a single cusp. This mathematical complexity mirrors the physical intricacy of the regime under consideration. By incorporating the coupling between hydrodynamic and diffusive-thermal effects, the proposed model offers a more robust framework for interpreting the fine-scale cellular structures that have been observed in experiments on flames near the threshold of stability. Within this distinguished crossover limit, the hydro-diffusive area $\mathcal{A}$ serves as a robust coupling constant. Whereas the Markstein number $\mathcal{M}$ is highly sensitive to the Lewis number, changing sign as the DT instability threshold is crossed, the coefficient $\mathcal{N}$ remains a stable, leading-order interaction term. To a first approximation, $\mathcal{N}$ may be treated as independent of the Lewis number in this narrow regime; however, it may be expected to vary as the internal flame structure undergoes more substantial modifications farther from the threshold. A first-principles derivation of the coefficient $\mathcal{N}$ from the full Navier–Stokes equations remains a challenging undertaking, requiring an asymptotic development that goes beyond the established Matalon–Matkowsky–Clavin–Joulin theory. Its determination may prove more tractable within the framework of Darcy’s law, where the hydrodynamic description is considerably simplified rajamanickam2024hydrodynamic, rajamanickam2026flame. Alternatively, $\mathcal{N}$ might be inferred from experimental measurements of cellular flame structures or extracted from direct numerical simulations, thereby providing a practical route to quantifying the hydro-diffusive area $\mathcal{A}$. More specifically, the limit $\mathcal{M}\to 0^{\pm}$ is of particular interest; in this limit, the classical theory, with its assumption of an order-unity Markstein number, becomes inapplicable. A new hydrodynamic description of premixed flames is therefore required, a direction that is both warranted and reserved for future investigation.
@@
Note that the present study has been conducted within the framework of a weak-heat-release approximation, adopted in the interest of simplicity. The qualitative features of the discussion may be expected to carry over, with appropriate modifications, when this approximation is relaxed. In the latter case, for instance, the quadratic stabilisation vanishes not simply when $\mathcal{M}\leq 0$, but rather when $\mathcal{M}-\mathcal{M}_{c}\leq 0$, where $\mathcal{M}_{c}$ denotes a critical Markstein number influenced by effects such as viscosity (see, for example, formula (19) and $\S$3.2 in matalon2018darrieus, which follows from a Taylor-series expansion of the Clavin–Garcia equation).
@@
Finally, it is appropriate to conclude this note by comparing (see Fig. 6) the distinguished DL-DT crossover regime proposed by Sivashinsky sivashinsky1988nonlinear, matalon2007intrinsic for small values of $\mathcal{M}$ (or $\mathcal{M}-\mathcal{M}_{c}$) with ours:
@@
or, alternatively,
Sivashinsky: $\quad\sigma=\varepsilon|k|-\mathcal{M}k^{2}-4k^{4}\quad$ $\Rightarrow\quad k\sim\sqrt{\mathcal{M}},\quad\sigma\sim\mathcal{M}^{2},\quad f\sim\mathcal{M},\quad\varepsilon\sim\mathcal{M}^{3/2},$
Current model: $\quad\sigma=\varepsilon|k|-\mathcal{M}k^{2}-\mathcal{N}|k|^{3},\quad$ $\Rightarrow\quad k\sim\mathcal{M},\quad\sigma\sim\mathcal{M}^{3},\quad f\sim\mathcal{M},\quad\varepsilon\sim\mathcal{M}^{2}.$
These scalings complement the pure DL regime scalings:
@@
The reader may also refer to a recent study creta2020propagation that explored nonlinear flame dynamics based on Sivashinsky’s scalings. Since our Markstein-number scaling $\mathcal{M}\sim\sqrt{\varepsilon}$ characteristic of the present model exceeds that of Sivashinsky $\mathcal{M}\sim\varepsilon^{2/3}$, the cubic stabilisation becomes operative earlier as the Markstein number is reduced toward the DT instability threshold. In other words, as a flame transitions from DL-modulated propagation toward the DT instability threshold, the cubic term engages first, providing the primary stabilising mechanism, while the quartic term only emerges in the deeper asymptotic limit where $\mathcal{M}$ is extremely small. Accordingly, when comparing cubic $(-|k|^{3})$ versus quartic $(-k^{4})$ stabilisation, the cubic model yields cellular structures of larger characteristic dimension, slower growth rates, and greater amplitudes. The present scaling thus captures the physically more immediate crossover regime where both hydrodynamic and diffusive-thermal instabilities participate on an equal footing.